OpenStudy (anonymous):

a resturaunt uses 7 1/2 pound of carrots to mak 5 carrot cakes. how many pounds of carrot are needed to make 1 cake???

OpenStudy (jdoe0001):

do the amounts as before \(\begin{array}{ccllll} lbs&cake \\\hline\\ 7\frac{1}{2}&5\\ x&1 \end{array}\implies \cfrac{7\frac{1}{2}}{x}=\cfrac{5}{1}\)

jhonyy9 (jhonyy9):

7 1/2 ----- 5 x ----- 1 ------------------ x = ?

OpenStudy (texaschic101):

7 1/2 divided by 5 15/2 * 1/5 = 15/10 = 1 1/2 lbs per cake

OpenStudy (anonymous):

THANK YOU SOO MUCH!!!!

OpenStudy (texaschic101):

yw